# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 6.3. SLM for Fine-tuning LLMs
SLMs as proxy models: SLMs can approximate the gradient of fine-tuning large-scale LLMs on target datasets, avoiding the costly fine-tuning process in terms of time and computational resources. As shown in Figure 23 (a), Emulated Fine-Tuning (EFT) (Mitchell et al., 2024) simulates both unsupervised pre-training and supervised fine-tuning stages across different scales by manipulating log probabilities. EFT, for example, combines base log probabilities from a 70B model with behavioral deltas from a 7B model—these deltas represent differences between fine-tuned and unfine-tuned SLMs, effectively emulating outcomes for the Llama-2 series. This method allows fine-tuning on smaller models such as Falcon-7B (Almazrouei et al., 2023) while capturing most benefits of fine-tuning larger models such as Falcon-180B, benefiting applications such as dialogue, question-answering, and code generation. Similarly, Proxy-tuning (Liu et al., 2024d) adjusts LLM predictions by adding the differences between the outputs of a fine-tuned small model and its untuned version to the LLM’s output vocabulary during decoding, maintaining the advantages of large-scale pre-training while integrating small-scale fine-tuning benefits.
SLMs play a role in selecting high-quality fine-tuning data for LLMs. Figure 23 (b) illustrates how SLMs within the same family as the LLM can identify training samples that are likely to be challenging, enhancing the training efficiency and generalization capability of the LLM. As demonstrated by Swayamdipta et al. (2020) and further advanced by Mekala et al. (2024), the learning percentage $LP(i)$ is a metric used to curate high-quality datasets with hard samples:
(10) $LP(i)=\frac{P_{i-1}-P_{i}}{P_{0}-P_{n}}$
where $P_{i}$ represents the perplexity at the end of epoch-$i$, and $P_{0}$ is the initial perplexity. A higher $LP(i)$ early in training indicates significant learning in the initial epochs, highlighting the potential of these samples to enhance LLMs.
@@ 6.4. SLM for LLM Applications
Figure 24. Contrastive Decoding (Li et al., 2023d).
@@
In open-ended text generation—such as writing assistance and story creation—LLMs often suffer from issues such as incoherence and thematic drift over extended sequences. Due to more frequent failure patterns observed in SLMs, such as short, repeated, and irrelevant strings, these patterns serve as negative examples for LLM decoding. Contrastive Decoding (CD) (Li et al., 2023d) improves coherence and lexical diversity by leveraging the differential capabilities between a large model, OPT-13B (Zhang et al., 2022), and a smaller model, OPT-125M. As illustrated in Figure 24, CD improves content quality by sampling generation based on the difference in log probabilities, $\log p_{EXP}-\log p_{AMA}$, between an expert LM and an amateur LM, rather than relying solely on the expert LM’s log probability. This approach effectively reduces generative failures, including repetition.
@@
In knowledge injection, general LLMs based on open-domain data may lack the domain-specific knowledge necessary for specialized tasks in fields such as law and medicine. In contrast, domain-specific SLMs can provide crucial domain knowledge and adapt this knowledge in a format that LLMs prefer. To address this, BLADE (Li et al., 2024a) augments black-box LLMs with small domain-specific models. The structure of BLADE comprises both the black-box LLMs, which offer strong language comprehension and reasoning capabilities, and small domain-specific LMs, which retain specialized insights. As illustrated in Figure 25, the BLADE approach involves three critical steps: 1) pre-training the SLM on domain-specific data to encapsulate the necessary expertise, 2) fine-tuning the model using knowledge instruction data to refine its application to task-specific requirements, and 3) employing joint Bayesian optimization to harmonize the capabilities of the general-purpose LLM with the small LM, enhancing overall performance.
@@
In relation extraction, a field limited by scarce labeled data and prevalent long-tail categories, the “Train-Guide-Predict” framework (Tang et al., 2024b) employs SLMs to learn task-specific knowledge for dominant categories. SLMs struggle with rare categories, whereas LLMs manage these effectively due to their extensive pre-trained text. Therefore, This framework leverages the strengths of both models: it utilizes SLMs for acquiring task knowledge and guiding the LLM’s generative process with initial SLM predictions, enhancing the LLM’s handling of underrepresented categories.
In generating empathetic responses, LLMs are proficient in expressive capabilities but often lack a nuanced understanding of emotions and cognition. Thus, the Hybrid Empathy Framework (HEF) (Yang et al., 2024d) enhances this by integrating Small Empathy Models (SEMs) to deepen the emotional and cognitive insight of LLMs. This framework uses a two-tiered emotion prediction approach: SEMs first determine primary emotions, which then guide the LLM to focus on these emotions and the key emotional triggers. This strategy leads to more accurate and empathetic responses.
@@ 6.5. SLM for LLM Evaluation
SLMs can also enhance the evaluation of LLMs, especially in dialogue generation, where handling diverse potential responses is inherently challenging. In dialog evaluation, generating both dialog and reference responses computationally is complex, making accurate assessment difficult due to the multiple plausible but semantically different responses possible for a single dialog context. Sole reliance on LLMs for evaluation can lead to problems such as dependency on prompt wording and inconsistent results. Therefore, a specifically trained SLM can address these issues. SLIDE framework (Zhao et al., 2024d) employs contrastive learning to fine-tune an SLM to effectively distinguish between positive and negative responses. This trained SLM is subsequently combined with an LLM to assign a score to each response, optimizing the evaluation process. The scoring method used is formalized as follows:
(11) $\small score=\begin{cases}score_{SLM},&\text{if }score_{SLM}\geq 0.5\\ score_{LLM},&\text{if }score_{LLM}<0.5\\ \frac{score_{SLM}+score_{LLM}}{2},&\text{otherwise}\end{cases}$
This equation allows for adaptive response evaluation, leveraging the strengths of both models to ensure a more reliable and consistent assessment across varying dialogue contexts.
@@ 7. Trustworthiness in Small Language Models
Language models have become ubiquitous in our daily lives, and we increasingly rely on them. However, they pose risks regarding their limitations in some trustworthy dimensions like privacy and fairness. These concerns are especially critical in high-stakes domains such as healthcare (He et al., 2023b) and finance (Li et al., 2023f). Consequently, numerous studies have emerged to evaluate the trustworthiness of LMs (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024). Though there are already a lot of works exploring LMs’ trustworthiness, most of them focus on LLMs. In this section, we investigate works that contain LMs of around 7B size and below. Note that we do not include the specific attack methods (Zou et al., 2023; Carlini et al., 2021; Chen et al., 2022b; Huang et al., 2023a) or work (Yang et al., 2023b) that only focuses on early pre-trained LMs like BERT (Devlin et al., 2019) as they are already covered in previous survey papers (Ramesh et al., 2023; Guo et al., 2022; Goyal et al., 2023; Delobelle et al., 2022). Inspired by previous works (Sun et al., 2024b; Wang et al., 2023a), we discuss the following five key trustworthy scenarios: robustness, privacy, reliability, safety, and fairness, as shown in Figure 26. We consider two scenarios for robustness: Adversarial (Adv) Robustness (Wang et al., 2021) and Out-of-Distribution (OOD) Robustness (Bulusu et al., 2020; Liu et al., 2021). For safety, we explore two key concerns: Misinformation (Van Der Linden, 2022) and Toxicity (Welbl et al., 2021). For reliability, we focus on Hallucination (Huang et al., 2023b) and Sycophancy (Sharma et al., 2023). Please note that these are just the aspects we are focusing on, and therefore this is not a comprehensive classification or taxonomy. For example, robustness also contains robustness to adversarial demonstration. Next, we briefly introduce some representative works.
@@
Do-Not-Answer (Wang et al., 2024d) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
@@
HaluEval (Li et al., 2023c) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
@@
PrivLM-Bench (Li et al., 2024b) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$

### v2

@@ 6.3. SLM for Fine-tuning LLMs
SLMs as proxy models: SLMs can approximate the gradient of fine-tuning large-scale LLMs on target datasets, avoiding the costly fine-tuning process in terms of time and computational resources. As shown in Figure 25 (a), Emulated Fine-Tuning (EFT) (Mitchell et al., 2024) simulates both unsupervised pre-training and supervised fine-tuning stages across different scales by manipulating log probabilities. EFT, for example, combines base log probabilities from a 70B model with behavioral deltas from a 7B model—these deltas represent differences between fine-tuned and unfine-tuned SLMs, effectively emulating outcomes for the Llama-2 series. This method allows fine-tuning on smaller models such as Falcon-7B (Almazrouei et al., 2023) while capturing most benefits of fine-tuning larger models such as Falcon-180B, benefiting applications such as dialogue, question-answering, and code generation. Similarly, Proxy-tuning (Liu et al., 2024d) adjusts LLM predictions by adding the differences between the outputs of a fine-tuned small model and its untuned version to the LLM’s output vocabulary during decoding, maintaining the advantages of large-scale pre-training while integrating small-scale fine-tuning benefits. Moreover, SLMs can act as proxies for approximate LLM fine-tuning during decoding. Weak-to-Strong Search (Zhou et al., 2024) strategy frames the alignment of LLMs as a test-time greedy search, aiming to maximize the log-probability difference between small tuned and untuned models while sampling from the frozen large model. This approach serves as a dual-purpose method: (1) a compute-efficient model up-scaling strategy that circumvents direct tuning of the large model, and (2) an instance of weak-to-strong generalization that bolsters a strong model with weak test-time guidance.
SLMs play a role in selecting high-quality fine-tuning data for LLMs. Figure 25 (b) illustrates how SLMs within the same family as the LLM can identify training samples that are likely to be challenging, enhancing the training efficiency and generalization capability of the LLM. As demonstrated by Swayamdipta et al. (2020) and further advanced by Mekala et al. (2024), the learning percentage $LP(i)$ is a metric used to curate high-quality datasets with hard samples: $LP(i)=\frac{P_{i-1}-P_{i}}{P_{0}-P_{n}}$ where $P_{i}$ represents the perplexity at the end of epoch-$i$, and $P_{0}$ is the initial perplexity. A higher $LP(i)$ early in training indicates significant learning in the initial epochs, highlighting the potential of these samples to enhance LLMs. SmallToLarge (S2L) (Yang et al., 2024d) utilizes training loss trajectories from smaller models to guide data selection for larger models fine-tuning. Experimental results demonstrate that S2L significantly enhances data efficiency in SFT for mathematical problem-solving, reducing the required training data to just 11% of the original MathInstruct dataset (Yue et al., [n. d.]) to achieve performance comparable to that obtained using the full dataset.
@@ 6.4. SLM for LLM Applications
Figure 26. Contrastive Decoding (Li et al., 2023e).
@@
In open-ended text generation—such as writing assistance and story creation—LLMs often suffer from issues such as incoherence and thematic drift over extended sequences. Due to more frequent failure patterns observed in SLMs, such as short, repeated, and irrelevant strings, these patterns serve as negative examples for LLM decoding. Contrastive Decoding (CD) (Li et al., 2023e) improves coherence and lexical diversity by leveraging the differential capabilities between a large model, OPT-13B (Zhang et al., 2022), and a smaller model, OPT-125M. As illustrated in Figure 26, CD improves content quality by sampling generation based on the difference in log probabilities, $\log p_{EXP}-\log p_{AMA}$, between an expert LM and an amateur LM, rather than relying solely on the expert LM’s log probability. This approach effectively reduces generative failures, including repetition.
@@
In knowledge injection, general LLMs may lack domain-specific expertise for specialized tasks like law or medicine (Wang et al., 2024a; Dou et al., 2023). Domain-specific SLMs can supply crucial knowledge in a format suitable for LLMs. To this end, BLADE (Li et al., 2024a) integrates black-box LLMs with small domain-specific models. BLADE combines the comprehensive language capabilities of LLMs with the specialized knowledge of small LMs. As shown in Figure 27, BLADE’s process includes: 1) pre-training SLMs on domain-specific data, 2) fine-tuning with knowledge instruction to meet task-specific needs, and 3) using joint Bayesian optimization to enhance synergy between the LLM and the small LM, boosting overall performance.
@@
In relation extraction, a field limited by scarce labeled data and prevalent long-tail categories, the “Train-Guide-Predict” framework (Tang et al., 2024b) employs SLMs to learn task-specific knowledge for dominant categories. SLMs struggle with rare categories, whereas LLMs manage these effectively due to their extensive pre-trained text. Therefore, this framework leverages the strengths of both models: it utilizes SLMs to acquire task knowledge and guide the LLM’s generative process with initial SLM predictions, enhancing the LLM’s handling of underrepresented categories.
In generating empathetic responses, LLMs excel in expressiveness but struggle with nuanced emotions and cognition. HEF (Yang et al., 2024e) addresses this by incorporating Small Empathy Models (SEMs) to enhance LLMs’ emotional and cognitive depth. This framework employs a two-tiered emotion prediction method: SEMs identify primary emotions, directing LLMs to concentrate on these emotions and their triggers, resulting in more accurate and empathetic responses.
6.5. SLM for LLM Safety
As demonstrated by various works (Shayegani et al., 2023; Zhuo et al., 2023; Mozes et al., 2023; Yuan et al., 2024a), LLMs are vulnerable to adversarial attacks and jailbreaking. For example, Wang et al. ([n. d.]) shows that ChatGPT’s performance on adversarial datasets is still far from perfect, indicating that potential risks of adversarial vulnerability remain. Another example includes jailbreaking ChatGPT by asking it to ’pretend to be a sarcastic mean girl.’ Using such techniques, it has been shown that even the most advanced LLMs are far from being safe against generating potentially harmful content. Hence, the widely adopted LLM-based services to generate are at high risk of being misused for nefarious purposes. Consequently, resources such as the Llama 2 Responsible Use Guide  [^14] https://ai.meta.com/static-resource/responsible-use-guide/ strongly advocate for implementing robust guardrails in products that utilize Generative AI. These guardrails are specifically designed to mitigate risks associated with both inputs to and outputs from the model, ensuring safeguards against the generation of high-risk or policy-violating content, as well as protecting against adversarial inputs and attempts to compromise the model. In addition to developing trustworthy LLMs, adopting SLMs for LLM safety (Inan et al., 2023; Kwon et al., 2024) has also attracted increasing attention. For example, Llama Guard (Inan et al., 2023), fine-tuned on Llama2-7B, has publicly released an input-output safeguard tool specifically for classifying safety risks in prompts and responses within conversational AI applications. However, this tool is limited to assessing the harmfulness of questions and answers and does not facilitate the generation of fluent, safe responses. In response to this limitation, Kwon et al. (2024) fine-tune a specialized small language model with harmful query detection and safeguard answer generation tasks to accurately detect harmful user queries and generate appropriate safeguard explanations, thereby enhancing the safety measures in conversational AI.
@@ 6.6. SLM for LLM Evaluation
SLMs can also enhance the evaluation of LLMs. In dialog evaluation, generating dialog reference responses is computationally complex, making accurate assessment difficult due to the multiple plausible but semantically different responses possible for a single dialog context. Relying on LLM prompting for evaluation can lead to problems such as dependency on prompt wording and inconsistent results. One solution involves training specialized SLMs to evaluate LLMs, as these SLMs can be fine-tuned more quickly and generate outputs faster during inference, owing to their reduced number of parameters. For example, SLIDE (Zhao et al., 2024d) employs contrastive learning to fine-tune an SLM to effectively distinguish between positive and negative responses. Based on its observation that SLMs are more accurate in identifying positive responses and LLMs excel at classifying negative ones, the trained SLM is subsequently integrated with an LLM to assign a score to each response. The scoring method used is formalized as follows:
(6) $\small score=\begin{cases}score_{SLM},&\text{if }score_{SLM}\geq 0.5\\ score_{LLM},&\text{elif }score_{LLM}<0.5\\ \frac{score_{SLM}+score_{LLM}}{2},&\text{otherwise}\end{cases}$
This equation allows for adaptive response evaluation, leveraging the strengths of both models to ensure a more reliable and consistent assessment across varying dialogue contexts. In the natural language generation task,  Kuhn et al. (2023) designs a novel entropy to evaluate the uncertainty of LLMs. It aims to tackle the challenge of semantic equivalence (Kuhn et al., 2023). For instance, A’s son is B and B is A’s son are semantically equivalent. It should not be considered uncertain if an LLM is unsure about which of the two previously mentioned sentences to generate due to semantic equivalence. A DeBERTa-Large (He et al., 2020) fine-tuned on the MNLI (Williams et al., 2018) dataset serves as the classifier guided by semantic equivalence in the clustering stage. SelfCheckGPT (Manakul et al., 2023) proposes a black-box hallucination detection method for LLMs. The core idea is to leverage uncertainty derived from sampled outputs. To be specific, Manakul et al. (2023) claim that an LLM trained on a concept generates responses that are similar and factually consistent. One of the five variants of SelfCheckGPT uses BertScore to achieve it. A DeBERTa-Large (Liu et al., 2019) is utilized to calculate the BERTScore. Factscore (Min et al., 2023) is proposed to evaluate the factuality of LM-generated long-form content. It divides the generated long content into multiple short texts, enabling a more precise assessment of factual accuracy. In addition to manual evaluation, Min et al. (2023) also proposes an automated evaluation framework to estimate Factscore which can reduce costs. LLaMa 7B (Touvron et al., 2023a) fine-tuned on Super-NaturalInstructions (Wang et al., 2022a) is one of the LMs employed as an evaluation assistant and shows promising performance. They also employ Generalizable T5-based dense retrievers (Ni et al., 2022) to facilitate passage retrieval.
@@
7. Synergy between Small and Large Language Models
The synergy between small and large language models leverages the unique strengths of each to enhance overall system performance and efficiency. SLMs, being lightweight and resource-efficient, are ideal for deployment on edge devices, enabling rapid responses and low latency for straightforward tasks. LLMs, on the other hand, possess greater computational power and a deeper understanding of complex language patterns, allowing them to handle more intricate and nuanced tasks. By integrating SLMs and LLMs, systems can dynamically allocate tasks based on complexity, ensuring that simple queries are processed quickly on the edge while more demanding requests are escalated to the cloud. This collaborative approach optimizes resource usage, reduces operational costs, and maintains high-quality outputs across a diverse range of applications. The synergy between small and large language models can be categorized into two parts: cloud-edge synergy and task-centric synergy. Cloud-edge synergy refers to a setup where SLMs operate on edge devices, while LLMs reside on the server. When the SLM is not powerful enough, the LLM compensates by handling more complex tasks and providing additional support. Task-centric synergy refers to the scenario where SLMs and LLMs leverage their respective strengths to improve task-oriented efficiency. Table 13 summarizes representative work in each category and their key points. Next, we introduce each category in detail.
Table 13. Synergy between SLMs and LLMs
Synergy Representative Work Key Point
Cloud-Edge Synergy (Inference) CoGenesis (Zhang et al., 2024h) Divide user instructions into general part by LLMs and private parts by SLMs.
Xu et al. (2024a) Introduce split learning in 6G to distribute LLM agents.
LLM-to-SLM (Bergner et al., 2024) Encode prompts with server-side LLM and decodes with edge-side SLM.
Synergy of Thoughts (Shang et al., 2024) SLMs suggest reasoning paths; LLMs correct contradictions.
Hao et al. (2024) SLM generates local tokens; LLM checks and corrects complex tokens.
LLMCad (Xu et al., 2023d) Combine lightweight and high-precision LLMs for on-device inference.
Khattab et al. (2023); Ma et al. (2023a) Focuse on LLM’s reasoning and SLM’s efficient decoding.
Cloud-Edge Synergy (Training) CROSSLM (Deng et al., 2023) Preserve client data privacy by training SLM locally and LLM remotely; mutual improvement using SLM-labeled data from LLM outputs.
Task-Centric Synergy $\alpha$-UMi (Shen et al., 2024a) Break down a single LLM into specialized agents.
SynCID (Liang et al., 2024) Merge LLM’s semantic with SLM’s speed; refine labels via contrastive learning.
Filter-then-rerank (Ma et al., 2023a) SLMs process simple samples and flag complex ones for LLM reranking.
Data Shunt+ (DS+) (Chen et al., 2024c) Process easy samples with SLMs and delegates hard samples to LLMs.
Figure 28. Could-edge synergy between LLMs and SLMs.
7.1. Cloud-Edge Synergy
The current utilization of LLMs typically involves uploading private data to the cloud for response. Fine-tuning LLMs usually also requires uploading data to clouds for computing. However, this raises privacy concerns as the collection and usage of private data are constrained by personal privacy awareness and legal regulations (Voigt and Von dem Bussche, 2017). Consequently, the cloud-edge synergy between SLMs and LLMs is proposed to alleviate this issue, i.e., SLMs handle privacy-sensitive data locally, LLMs handle de-identified or non-privacy-sensitive data, and these two models collaborate. This section discusses such cloud-edge synergy, dividing them into two categories: cloud-edge synergy during inference and cloud-edge synergy during training, as shown in Figure 28.
Cloud-edge Synergy During Inference. CoGenesis (Zhang et al., 2024h) breaks down the user instruction into a general section and a personal section. The LLM generates replies solely based on general instruction, and the SLM considers both user instruction and additional personal context for its output generation. A fusion strategy blends the output of LLM and SLM synergistically. Xu et al. (2024a) introduces a split learning system for LLM agents in 6G networks, optimizing mobile device and cloud server collaboration. Mobile devices operate lightweight SLMs with 0–10B parameters for real-time tasks, while cloud servers handle larger LLMs with over 10B parameters for complex reasoning and planning. This setup allows efficient local task management on mobile devices and offloads heavy operations to cloud servers. The system’s architecture features three modules—perception, grounding, and alignment—facilitating effective communication to meet the sophisticated needs of 6G networks.
Besides these frameworks, more specific models are proposed to facilitate the cloud-edge synergy. A common strategy is to use SLM’s fast decoding ability. LLM-to-SLM (Bergner et al., 2024) proposes a framework in which the pre-trained frozen encoder-decoder LLM resides on the server and computes a high-quality representation of the prompt for the planning of an appropriate response. The SLM residing on the edge device, conditioned on this representation, decodes the response efficiently. Some variants put more emphasis on the reasoning ability of LLMs (Khattab et al., 2023; Shang et al., 2024; Ma et al., 2023a). In Synergy of Thoughts (Shang et al., 2024), the SLMs generate multiple low-cost reasoning paths. If these paths conflict, the larger LLMs are invoked to provide reflective reasoning and correct any intuitive errors. Hao et al. (2024) proposes a framework in which an SLM residing on the edge devices generates tokens, calling LLMs to verify and correct threshold-gated ”harder” tokens, to achieve a controllable trade-off between inference quality and cost. LLMCad (Xu et al., 2023d) presents an on-device inference engine addressing memory and latency issues in deploying LLMs on mobile devices. It combines a lightweight LM for token generation with a high-precision LLM for verification, leveraging a token tree structure and speculative generation for efficiency. Tested on devices such as Jetson TX2, it achieves up to 9.3× speedup for LLMs with over 10 billion parameters while maintaining accuracy.
Cloud-edge Synergy During Training. CROSSLM (Deng et al., 2023) introduces a client-server collaborative training framework that preserves data privacy by having clients locally train SLMs instead of fine-tuning LLMs. The framework enables mutual enhancement through a feedback loop where SLMs evaluate LLM-generated synthetic data and provide feedback to improve the LLM’s generative capabilities, ensuring high-quality and task-specific data. Concurrently, the synthetic data trains the SLMs, boosting their performance. This cyclical exchange fosters cloud-edge synergy and mutual model improvement.
7.2. Task-Centric Synergy
The advent of LLMs has significantly propelled various natural language processing tasks and inspired research into their synergistic interactions with SLMs to enhance the performance of models tailored for specific tasks. This section introduces scenarios where small language models exhibit specialized capabilities after fine-tuning and discusses how combining their unique strengths with the versatile abilities of LLMs can yield superior performance on specific tasks. For example, LLMs excel at handling difficult examples or can rewrite content to eliminate task-irrelevant redundancy, thereby enhancing overall task performance, as illustrated in Figure 29, 30 and 31.
Figure 29. Synergizing SLMs and LLMs in tool learning.
$\alpha$-UMi (Shen et al., 2024a) introduces a multi-agent framework to enhance tool learning by overcoming the limitations of single-LLM approaches for complex tasks. It utilizes three specialized LMs—planner, caller, and summarizer—as depicted in Figure 29—each handling specific subtasks such as planning, tool invocation, and summarization. This modular design allows the use of small and large open-source LLMs (e.g., LLaMa-7B/12B) and supports easy tool updates. Evaluated on benchmarks like ToolBench (Qin et al., [n. d.]) and ToolAlpaca (Tang et al., 2023), $\alpha$-UMi outperforms traditional single-LLM methods and even exceeds GPT-4 in tool learning performance.
Figure 30. Synergizing SLMs and LLMs in Conversational Intent Detection.
SynCID (Liang et al., 2024) focuses on Conversational Intent Discovery (CID), a task where both known and new intents must be identified from user utterances in an open-world setting. SynCID combines LLMs’ deep semantic insights with SLMs’ agility and specialized capabilities. As illustrated in Figure 30, the framework uses LLM prompting to refine discourse and intent labels, enhancing semantic accuracy and assigning new labels to unlabeled data. SLMs are trained via contrastive learning to align semantic spaces of discourse and intent descriptors, reducing clustering distortion and improving new intent detection. Tested on BANKING (Casanueva et al., 2020), CLINC (Larson et al., 2019), and StackOverflow (Xu et al., 2015), SynCID outperforms CID baselines significantly.
Figure 31. Synergizing SLMs and LLMs in Information Extraction.
Filter-then-rerank (Ma et al., 2023a) addresses LLMs’ poor performance on simpler IE tasks by integrating LLMs and SLMs. SLMs act as filters, predicting and identifying difficult samples, while LLMs rerank the top N candidate labels for these cases. As illustrated in Figure 31, SLM predictions are final for non-difficult samples, minimizing reliance on LLMs and reducing latency and costs; for those difficult samples, the top N predicted candidate labels from the SLM are passed to the LLM for reranking (predicting). Tested on small-sample IE tasks, this approach improves performance by an average of 2.4% compared to previous methods. Data Shunt+ (DS+) (Chen et al., 2024c) introduces a framework to reduce costs by minimizing large model queries during inference and boosting LLM performance with SLMs for tasks like sentiment analysis and image processing. DS+ uses SLMs for “easy” samples within the main training distribution and LLMs for ”hard” outliers or boundary cases, maintaining accuracy while reducing LLM use. It incorporates S4L and L4S modules with Prompt Pruning (PP) and 2-stage Confidence Distillation (2CD) for better input processing and knowledge transfer. Tests show DS+ outperforms fine-tuning in accuracy and cost efficiency, significantly cutting down on LLM queries.
@@ 8. Trustworthiness in Small Language Models
Language models have become ubiquitous in our daily lives, and we increasingly rely on them. However, they pose risks regarding their limitations in trustworthy dimensions like privacy and fairness. These concerns are especially critical in high-stakes domains such as healthcare (He et al., 2023b) and finance (Li et al., 2023g). Consequently, numerous studies have emerged to evaluate the trustworthiness of LMs (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024b). In this section, we consider the works that benchmark various LMs’ trustworthiness and omit the specific attack methods (Zou et al., 2023; Carlini et al., 2021; Chen et al., 2022b; Huang et al., 2023a) or work (Yang et al., 2023b) that only focuses on early pre-trained LMs like BERT (Devlin et al., 2019) as they are already covered in previous survey papers (Ramesh et al., 2023; Guo et al., 2022; Goyal et al., 2023; Delobelle et al., 2022). Inspired by previous works (Sun et al., 2024b; Wang et al., 2023a), we discuss the following five key trustworthy scenarios: robustness, privacy, reliability, safety, and fairness, as shown in Figure 32. We consider two dimensions for robustness: Adversarial (Adv) Robustness (Wang et al., 2021b) and Out-of-Distribution (OOD) Robustness (Bulusu et al., 2020; Liu et al., 2021). For safety, we explore two key concerns: Misinformation (Van Der Linden, 2022) and Toxicity (Welbl et al., 2021). For reliability, we focus on Hallucination (Huang et al., 2023c) and Sycophancy (Sharma et al., 2023). Please note that these are just the aspects we are focusing on, and therefore this is not a comprehensive classification or taxonomy. For example, robustness also contains robustness to adversarial demonstration.
Though there are a lot of works benchmarking LMs’ trustworthiness, their main focus is on LLMs. Therefore, we survey some representative works evaluating the trustworthiness of LMs, focusing specifically on those that include SLMs of around 7B parameters or smaller. We also summarize these works in Table 14. Next, we briefly introduce them.
@@
Do-Not-Answer (Wang et al., 2024f) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
@@
HaluEval (Li et al., 2023d) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
@@
PrivLM-Bench (Li et al., 2024c) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$